OpenStudy (anonymous):

HELP ME PLEASE :( The measures of the angles of a triangle are 2x, 3x, and 4x. What is the degree measure of the largest angle of the triangle? A.) 20° B.) 60° C.) 80° D.)100°

OpenStudy (texaschic101):

the angles of a triangle add up to 180 degrees so your problem is : 2x + 3x + 4x = 180 9x = 180 x = 20 2x = 2(20) = 40 degrees 3x = 3(20) = 60 degrees 4x = 4(20) = 80 degrees 40 + 60 + 80 = 180 so 4x (the largest angle) would be 4(20) = 80 degrees

OpenStudy (anonymous):

Thank you so much!

OpenStudy (texaschic101):

anytime :)